OpenStudy (jennychan12):

How to do Errors in Trapezoidal Rule and Simpson's Rule? See below for question.

OpenStudy (jennychan12):

\[\int\limits_{0}^{2}x^3dx\] n=4

OpenStudy (jennychan12):

I know the formula is \[E \le \frac{ (b-a)^2 }{ 12n^2 } (\max \left| f"x \right|), a \le x \le b\] and I got 1/4 but the answer is 1/2 am i doing something wrong?

OpenStudy (jennychan12):

@Hero @zepdrix

OpenStudy (jennychan12):

someone please help????

OpenStudy (jennychan12):

ok, i got the answer, but i don't get what they did in the last step. (maximum when x=2?)